# Mammo-Clustering:A Weakly Supervised Multi-view Global-Local Context Clustering Network for Detection and Classification in Mammography

### v1

@@ II-B2 Context Cluster Block
where $P^{x}$ represents a $x$ dimension of the central anchor point, $x\in\{r,g,b,h,w\}$. And $p_{q}^{x}$ represents the $n$-th neighbor point in the $x$ dimension, $q\in\{0,1,2,\ldots,k\}$.
After computing the features for all central anchor points, a similarity analysis is conducted between all points in the point set and each central anchor point’s features in step d. Each point is assigned to the cluster of the central anchor point with which it has the highest similarity. The steps for the similarity analysis is conducted by computing the pairwise cosine similarity matrix between a point and set of central points:
$f(P_{i},P^{\prime}_{j})=\frac{P_{i}\cdot P^{\prime}_{j}}{|P_{i}||P^{\prime}_{j}|}$
where $f(P_{i},P^{\prime}_{j})$ is pairwise cosine similarity compute. $P_{i}$ is $i$-th central anchor point, $i\in\{0,1,2,\ldots,n\}$. $P^{\prime}_{j}$ is $j$-th point in the image point set.
@@
Since each point contains both feature and positional information, the calculation implicitly emphasizes the point’s distance (locality) and feature similarity. Each point is then assigned to the most similar center, resulting in c clusters. During this process, clusters may contain varying numbers of points. In extreme cases, some clusters may contain zero points, rendering them redundant.
II-B3 ROI Selection Module
Firstly, the dimensions of the region of interest mapping are required to be consistent with the dimensions of the saliency-map.
$h_{crop}=h_{patch}\times h_{map}\div h_{I}$
$w_{crop}=w_{patch}\times w_{map}\div w_{I}$
where $h_{patch}$ and $w_{patch}$ are user-defined for the patch image size. $h_{map}$ and $w_{map}$ are the size of the saliency-map. $h_{I}$ and $w_{I}$ are the size of the original image data. $h_{crop}$ and $w_{crop}$ are dynamically adjusted based on the original image and saliency map scales. Subsequently, the saliency-map is normalized and divided into regions with a height of $h_{crop}$ and a width of $w_{crop}$ regions for greedy ROI search. In each iteration, the algorithm greedily identifies each region and selects the one with the largest total weight, as determined by average pooling, among all current regions. The coordinates of this region are added to a list, and a mask flag is applied to the region to prevent redundant selection. The coordinates of these regions will be mapped to the size of the original image to obtain patch-based images.
@@ II-B3 ROI Selection Module
Single-View Res[37] $0.727\pm 0.02$ 0.783 0.619 $0.719\pm 0.02$ 0.591 0.558 1477025
Single-View Swin-transformer $0.731\pm 0.02$ 0.651 0.594 $0.724\pm 0.02$ 0.601 0.599 14184625
GMIC[27] $0.793\pm 0.02$ 0.624 0.518 $0.778\pm 0.02$ 0.712 0.705 22487298
Multi-View res[37] $0.740\pm 0.02$ 0.753 0.567 $0.731\pm 0.02$ 0.676 0.630 6128546
MaMVT[38] $0.770\pm 0.02$ 0.918 0.647 $0.749\pm 0.02$ 0.649 0.649 30730082
Multi-View GMIC[52] $0.797\pm 0.02$ 0.637 0.521 $0.781\pm 0.02$ 0.719 0.699 22686871
Mammo-Clustering(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.694 $\textbf{0.805}\pm 0.02$ 0.709 0.709 9805459
TABLE I: Performance of each model on two datasets
II-B4 Attention Module
In our model, we use two different attention mechanisms to separately fuse multi-instance and multi-view information.
Similar Multi-instances information fusion: After the global network, the ROI select module to choose $k$ patch-based images, a number set manually. This implies not all patch-based images carry beneficial information, and some may be redundant. Considering and integrating all the information from these patch-based images could significantly impair our network. Therefore, an attention module is added before integrating local and global information, allowing the model to learn how to filter out irrelevant local information.
@@ II-B4 Attention Module
First, we use a neural network layer with simple linear transformations $f_{weights}$ and softmax function to compute attention weights $W_{att}$.
$W_{att}=\text{softmax}(f_{weights}(F_{l}))$
$F_{a}=F_{l}\odot W_{att}$
where $\odot$ represents the stationary point multiplication algorithm. Subsequently, the attention weights $W_{att}$ are multiplied pointwise with the feature representation of patch-based images $F_{l}$ is performed to obtain the final implicit representation $F_{a}$.
Multi-view information fusion: In multi-view learning, not all information from each view is necessarily classified as malignant. However, If a single view exhibits malignant characteristics, the instance should be classified as malignant. Therefore, we introduce an attention mechanism to enable the model to autonomously filter out irrelevant view information, enhancing classification accuracy.
The attention mechanism processing is largely consistent with multi-instance fusion attention. However, in multi-view attention, this attention module processes not only the $F_{a}$ fused by the multi-instance attention module but also $F_{g}$ and $F_{l}$. This is because all three features are considered in the loss function for loss calculation.
II-B5 Embedding Module
The embedding module in the model primarily aligns feature-based local information $F_{fl}$ with patch-based local information $F_{pl}$ before their integration. Here, we employ a trainable MLP to align the scales. We designed relevant ablation experiments to verify its effectiveness.
II-B6 Maxpooling Module
@@ II-C Loss Function
$L_{total}=\alpha\cdot L_{global}+\beta\cdot L_{local}+\gamma\cdot L_{fusion}+\delta\cdot L_{map}$
@@
$L_{global}$ is calculated using the global information obtained from multi-view fusion and the ground truth values. The loss function chosen here is BCELoss. And $L_{map}$ will be calculated from the saliency-map, it is the weighted average intensity of the saliency-map under the L1 norm. The $L_{global}$, combined with $L_{map}$, indicates the quality of the Global Network and further to determining the adjustment magnitude for the Global Network to enhance the accuracy of locating patch-based images. BCEWithLogitsLoss function is used for both $L_{local}$ and $L_{fusion}$. $L_{local}$ represents the quality of the local network, calculated from local information and ground truth values, determining the adjustment magnitude for the Local Network to enhance the feature extraction capability of the Local Network. $L_{fusion}$ represents the model’s final classification error, driving the overall model training. The weights $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the proportion of each loss, all manually set.
Figure 3: Figures a to d provide a visualization of patch-based images extracted by the model. The green box on the mammography indicates the location of the suspicious lesion, while the blue box represents the patch-based images selected by the model. We can observe that the model’s extracted patch-based images perform exceptionally well, and the magnified images clearly show calcifications and masses.
@@ III-B Evaluating Indicator
III-B1 AUC (Area Under the Curve)
AUC means the area under the receiver operating characteristic (ROC) curve. The ROC curve uses the true positive rate for mammography benign-malignant classification as the y-axis and the false positive rate as the x-axis. It provides an aggregate measure of performance across all possible classification thresholds. A higher AUC value indicates a better model performance, with 1 representing a perfect model and 0.5 a random guess.
@@ III-B1 AUC (Area Under the Curve)
where $TPR(t)$ is the true positive rate at threshold t, and $FPR(t)$ is the false positive rate at threshold t
III-B2 ACC (Accuracy)
Accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined. It gives a straightforward measure of how often the classifier is correct.
@@ III-B2 ACC (Accuracy)
III-B3 F1 Score
The F1 Score is the weighted average of Precision and Recall. This score takes both false positives and false negatives into account. It is particularly useful when the class distribution is uneven.
@@ III-B3 F1 Score
The F1 score reaches its best value at 1 (perfect precision and recall) and worst at 0.
@@ III-C Comparative Experiment
In this study, we evaluated several models on two datasets: Vindr-Mammo and cbis-ddsm. The performance metrics considered were AUC, ACC, and F1 score.
III-C1 Vindr-Mammo
For the Vindr-mammo, ours model achieved the highest performance across all metrics, with an AUC of $0.828\pm 0.02$, ACC of $0.919$, and F1 score of $0.694$. For single-view task models, the GMIC with a Global-local structure achieves an AUC of 0.793, significantly outperforming ResNet and Swin-Transformer, which had AUCs of $0.727\pm 0.02$ and $0.731\pm 0.02$, respectively. For multi-view task models, our model achieves an AUC of 0.828, significantly surpassing other models. Multi-View GMIC also showed competitive performance with an AUC of $0.797\pm 0.02$, validating the effectiveness of the Global-local architecture, but its ACC and F1 score were lower at $0.637$ and $0.512$.
III-C2 cbis-ddsm
On the cbis-ddsm dataset, the ours model again demonstrated superior performance with an AUC of $0.805\pm 0.02$, ACC of $0.709$, and F1 score of $0.709$. Multi-View GMIC, which had an AUC of $0.781\pm 0.02$ and an F1 score of $0.699$, obtained the highest ACC on this dataset. In this dataset, the advantages of the Global-Local architecture are more pronounced, with ResNet and Swin-Transformer showing significant disadvantages in AUC.
III-C3 Model Complexity
@@
III-C4 ROC curve
Figure 4: Figures a and b compare the ROC curves of our model with other Single-view and Multi-view architectures on the Vindr-mammo dataset. Figures c and d present the ROC curves comparison on the CBIS-DDSM dataset.
@@ III-D Ablation Experiment
III-D1 Different Information Fusion Method
This ablation study aims to demonstrate the effectiveness of our proposed weakly supervised architecture on mammography.
@@ III-D1 Different Information Fusion Method
Single-view Coc $0.762\pm 0.02$ 0.711 0.627
Multi-View Coc $0.783\pm 0.02$ 0.815 0.658
Ours(mv+gl) $\textbf{0.828}\pm 0.02$ 0.919 0.694
TABLE III: Performance of different information fusion methods on the Vindr-Mammo
The table clearly demonstrates the superiority of our architecture, achieving the highest AUC as well as optimal ACC and F1 scores, indicating its balance in mammography tasks. $mv$ represents a multi-view learning approach, and $gl$ refers to our proposed weakly supervised framework.
III-D2 Different Local Information
@@
This ablation study aims to verify the effectiveness of the mechanism integrating feature-based local information with patch-based local information.
@@
patch-based local information $0.810\pm 0.02$ 0.873 0.678
feature-based local information $0.806\pm 0.02$ 0.895 0.659
ours $\textbf{0.828}\pm 0.02$ 0.919 0.694
TABLE IV: Performance of different local information on the Vindr-Mammo
We found that focusing on only one type of local information doesn’t yield better results. The AUC for patch-based local information is 0.810, and for feature-based local information, it’s 0.806. However, combining both achieved the best result, with an AUC of 0.828.

### v2

@@ II-B2 Context Clustering
where $P^{x}$ denotes the feature of the central anchor point in the $x$ dimension, $x\in\{r,g,b,h,w\}$. Meanwhile, $p_{q}^{x}$ represents the feature of the $q$-th neighboring point in the $x$ dimension, $q\in\{0,1,2,\ldots,k\}$.
After computing the features for all central anchor points, a similarity analysis is conducted between all points in the point set and each central anchor point’s features in step d. Each point is assigned to the cluster of the central anchor point with which it has the highest similarity. The steps for the similarity analysis is conducted by computing the pairwise cosine similarity matrix $M$ between a point and set of central points:
$M(P_{i},P^{\prime}_{j})=\frac{P_{i}\cdot P^{\prime}_{j}}{|P_{i}||P^{\prime}_{j}|}$
where $M(P_{i},P^{\prime}_{j})$ is pairwise cosine similarity matrix. $P_{i}$ is $i$-th central anchor point, $i\in\{0,1,2,\ldots,n\}$. $P^{\prime}_{j}$ is $j$-th point in the image point set.
@@
Since each point in the set includes color and position information, this grouping implicitly incorporates spatial and image information [17].
II-B3 CAM
Feature visualization is computed by a feedforward network to generate saliency maps. This CAM module consists of convolution layers with a kernel size of 1 and is integrated into the Global Network to enable iterative optimization.
II-B4 ROI Selection Module
The saliency map is normalized and divided into regions with a height of $h_{crop}$ and a width of $w_{crop}$ regions for greedy ROI search, $h_{crop}$ and $w_{crop}$ are dynamically adjusted based on the original image and saliency map scales. In each iteration, the algorithm greedily identifies each region and selects the one with the largest total weight, as determined by average pooling, among all current regions. The coordinates of this region are added to a list, and a mask flag is applied to the region to prevent redundant selection. The coordinates of these regions will be mapped to the size of the original image to obtain patch-based images.
@@ II-B4 ROI Selection Module
TABLE II: Performance of each model on two datasets. SV and MV represent Single-View and Multi-Views, respectively. Backbone of the network after "-".
@@
DsMIL[18] $0.605\pm 0.02$ 0.730 0.781 $0.697\pm 0.02$ 0.500 0.583 202116
AbMIL[18] $0.618\pm 0.02$ 0.776 0.825 $0.726\pm 0.02$ 0.571 0.671 396547
TransMIL[20] $0.631\pm 0.02$ 0.888 0.890 $0.739\pm 0.02$ 0.635 0.637 2544402
SV Res18 [9] $0.727\pm 0.02$ 0.783 0.821 $0.719\pm 0.02$ 0.646 0.639 1477025
SV SwinT[21] $0.731\pm 0.02$ 0.651 0.719 $0.724\pm 0.02$ 0.651 0.599 14184625
GMIC-Res18 [15] $0.793\pm 0.02$ 0.847 0.878 $0.778\pm 0.02$ 0.682 0.680 22487298
MV Res18 [9] $0.740\pm 0.02$ 0.753 0.796 $0.731\pm 0.02$ 0.676 0.641 6128546
MaMVT [10] $0.770\pm 0.02$ 0.882 0.867 $0.749\pm 0.02$ 0.649 0.649 30730082
MV GMIC-Res18 [16] $0.797\pm 0.02$ 0.887 0.879 $0.781\pm 0.02$ 0.699 0.691 22686871
MV GMIC-SwinT $0.799\pm 0.02$ 0.874 0.854 $0.785\pm 0.02$ 0.694 0.694 28873234
Mammo-Clustering(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.906 $\textbf{0.805}\pm 0.02$ 0.709 0.709 9805459
II-B5 Attention Module
In our model, we employ attention mechanisms in two locations to integrate multi-instance and multi-view information, respectively.
Approximate Multi-Instances information fusion: Attention is employed here to mitigate the impact of redundant image patches on training. After the global network, the ROI select module to choose $n$ patch-based images, a number set manually. This implies not all patch-based images carry beneficial information, and some may be redundant. Considering and integrating all the information from these patch-based images could significantly impair our network. Therefore, an attention module is added before integrating local and global information, allowing the model to learn how to filter out irrelevant local information.
@@ II-B5 Attention Module
Firstly, we use a neural network layer with a simple linear transformation $f_{weights}$ and a softmax function to compute attention weights. Subsequently, the attention weights are multiplied pointwise with the feature representation of the block-based image $F_{l}$ to obtain the final implicit representation $F_{a}$.
$F_{a}=F_{l}\odot\text{softmax}(f_{weights}(F_{l}))$
where $\odot$ represents the stationary point multiplication algorithm.
Multi-view information fusion: Attention is employed here to mitigate the impact of redundant image views on training.
In multi-view learning, not all information from each view is necessarily classified as malignant. However, If a single view exhibits malignant characteristics, the instance should be classified as malignant. Therefore, we introduce an attention mechanism to enable the model to autonomously filter out irrelevant view information, enhancing classification accuracy.
The attention mechanism processing is largely consistent with multi-instance fusion attention. However, in multi-view attention, this attention module processes not only the $F_{a}$ fused by the multi-instance attention module but also $F_{g}$ and $F_{l}$. This is because we will perform targeted optimization on different network structures based on the losses obtained from various features.
II-B6 Embedding Module
The embedding module in the model primarily aligns feature-based local information $F_{fl}$ with patch-based local information $F_{pl}$ before their integration. Here, we employ a trainable MLP to align the scales. We designed relevant ablation experiments to verify its effectiveness, the effectiveness can be observed in Table VI.
II-B7 Maxpooling Module
@@ II-C Loss Function
$L=\alpha\cdot L_{global}+\beta\cdot L_{local}+\gamma\cdot L_{fusion}+\delta\cdot L_{map}$
@@
$L_{global}$ is calculated using the global information obtained from multi-view fusion and the ground truth values. The loss function chosen here is BCELoss. And $L_{map}$ will be calculated from the saliency-map, it is the weighted average intensity of the saliency-map under the L1 norm. The $L_{global}$, combined with $L_{map}$, indicates the quality of the Global Network and further to determining the adjustment magnitude for the Global Network to enhance the accuracy of locating patch-based images. BCEWithLogitsLoss function is used for both $L_{local}$ and $L_{fusion}$. $L_{local}$ represents the quality of the local network, calculated from local information and ground truth values, determining the adjustment magnitude for the Local Network to enhance the feature extraction capability of the Local Network. $L_{fusion}$ represents the model’s final classification error, driving the overall model training. The weights $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the proportion of each loss, all manually set, Here, we consider $L_{global}$, $L_{local}$, and $L_{fusion}$ to have equal importance, so we recommend setting $\alpha$, $\beta$, and $\gamma$ uniformly to 1. Meanwhile, $L_{map}$ represents the weighted average intensity of the saliency map calculated under the L1 norm, resulting in a relatively large value. Therefore, we suggest setting $\delta$ to a value less than 0.001.
Figure 4: Visualization of patch-based images extracted by the model. The green box on the mammography indicates the location of the suspicious lesion, while the blue box represents the patch-based images selected by the model. We can observe that the model’s extracted patch-based images perform exceptionally well, and the magnified images clearly show calcifications and masses.
@@ III-A2 CBIS-DDSM
Both Vindr and CBIS-DDSM provide detailed annotations of lesion locations, but such annotation tasks are generally high-cost. By employing weakly supervised learning to enable the network to autonomously localize lesion positions, the cost of dataset creation can be significantly reduced.
TABLE III: The composition of data for the two datasets.
@@ III-B Evaluating Indicator
III-B1 Classification indicators
  • AUC (Area Under the Curve): AUC means the area under the receiver operating characteristic (ROC) curve. The ROC curve uses the true positive rate for mammography benign-malignant classification as the y-axis and the false positive rate as the x-axis. It provides an aggregate measure of performance across all possible classification thresholds. A higher AUC value indicates a better model performance, with 1 representing a perfect model and 0.5 a random guess.
@@ III-B1 Classification indicators
    where $TPR$ is the true positive rate, and $FPR$ is the false positive rate.
  • ACC (Accuracy): Accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined, the corresponding clinical term is “specificity”. It gives a straightforward measure of how often the classifier is correct.
@@
  • F1 Score: The F1 Score is the weighted average of Precision and Recall. This score takes both false positives and false negatives into account. Given the long-tail issue in the data, we selected the micro F1 score as the evaluation metric, it is particularly useful when the class distribution is uneven.
@@
III-B2 Localization indicators
  • MDR(Miss Detection Rate): MDR is defined as the percentage of the number of undetected suspicious lesion areas $N_{miss}$ relative to the total number of suspicious lesion areas $N_{gt}$. Because, in clinical practice, we are more concerned about lesions being undetected, i.e., false negatives, rather than false positives.
    $\textit{MDR}=\frac{N_{miss}}{N_{gt}}$
  • Recall: Recall is a metric used in object detection to evaluate a model’s ability to identify all relevant objects in an image. It measures the proportion of actual positive instances (i.e., objects that should be detected) correctly identified by the model. In this context, it reflects the model’s capability to detect all existing lesions.
    $Recall=\frac{TP}{TP+FN}$
@@ III-C Comparative Experiment
In this study, we evaluated several models on two datasets: Vindr-Mammo and cbis-ddsm. The performance metrics considered were AUC, ACC, and F1 score and so on.
III-C1 Assessment of classfication
For the Vindr-mammo, ours model achieved the highest performance across all metrics, with an AUC of $0.828\pm 0.02$, ACC of $0.919$, and F1 score of $0.906$. For Single-View task models, the GMIC-Res18 with a Global-local structure achieves an AUC of 0.793, significantly outperforming other Single-View model, which SV Res18 and SV SwinT had AUCs of $0.727\pm 0.02$ and $0.731\pm 0.02$, respectively. For Multi-views task models, our model achieves an AUC of 0.828, significantly surpassing other models. Multi-views GMIC-SwinT also showed competitive performance with an AUC of $0.799\pm 0.02$, but its ACC and F1 score were lower at $0.874$ and $0.854$.
On the CBIS-DDSM dataset, the ours model again demonstrated superior performance with an AUC of $0.805\pm 0.02$, ACC of $0.709$, and F1 score of $0.709$, obtained the highest performance across on this dataset. Multi-View GMIC-Res18, which had an AUC of $0.781\pm 0.02$ and an F1 score of $0.699$. Noticeably, the Single-View MIL method shows promising potential on this dataset, with TransMIL achieving an AUC of $0.739\pm 0.02$. This patch-based information approach significantly outperforms methods that focus solely on global information.
In this this two datasets, the advantages of Context Clustering and Tri-level Information Fusion architecture are more pronounced than other information fusion method, showing significant advantages in AUC.
III-C2 Model Complexity
@@
III-C3 ROC curve
Figure 5: Comparison of ROC curves of different models on two public datasets. Figures a and b compare the ROC curves of our model with other Single-view and Multi-view architectures on the Vindr-mammo dataset. Figures c and d present the ROC curves comparison on the CBIS-DDSM dataset.
@@ III-C3 ROC curve
III-C4 Assessment of localization
We compared the performance of three weakly supervised lesion localization models without comparing them to fully supervised models. Our aim was to validate the feasibility of training weakly supervised models on datasets without annotated lesion regions, which would significantly reduce the burden of creating mammography datasets.
TABLE IV: Comparison of Different Weakly Supervised Models in Lesion Localization Tasks.
MDR Recall
GMIC-Res18 0.433 0.476
MV GMIC-Res18 0.336 0.643
MV GMIC-Res34 0.412 0.479
MV GMIC-SwinT 0.322 0.650
Ours 0.294 0.685
From Table IV, we observe that our model achieves the lowest missed detection rate (MDR) of 0.294 and the lowest recall rate of 0.476, demonstrating its potential. All the comparative models employ the same weakly supervised lesion localization approach as our model, enabling lesion location identification using only classification labels, making them suitable for comparison. We selected four different models encompassing both CNN and attention mechanism paradigms and conducted evaluation analysis on two metrics: MDR and recall. Compared to the second-best model, our approach shows an improvement of approximately 0.03 in both MDR and recall.
III-C5 Visual comparison
Figure 6: Comparison of activation map visualizations from different feature extraction paradigms. The green boxes on the original images indicate the annotated true locations of the lesions.
We visualized the feature activation maps of different feature paradigms, as shown in Figure 6. In this comparison, we selected representative models, ResNet and ViT, as exemplars of CNN and attention mechanisms, respectively, to compare with the Context Clustering paradigm.
It is evident that the activation maps based on CNN perform the worst. although they sometimes identify the locations of regions of interest, they exhibit low contrast. Activation maps based on attention mechanisms and clustering both achieve relatively clear and accurate localization of regions of interest. However, the activation maps based on Context Clustering are cleaner, indicating they are subject to less interference.
@@ III-D Ablation Experiment
III-D1 Different Feature Extraction Paradigm
This ablation study aims to demonstrate the superiority of Context Clustering in feature extraction performance for Mammography through numerical analysis.
TABLE V: Performance of different feature extraction paradigm on the Vindr-Mammo
@@ III-D1 Different Feature Extraction Paradigm
SV Res18(CNN-base) [9] $0.727\pm 0.02$ 0.783 0.821
SV SwinT(Atten-base) [21] $0.731\pm 0.02$ 0.651 0.719
SV Coc(Clustering-base) [17] $\textbf{0.762}\pm 0.02$ 0.794 0.833
The table clearly demonstrates the superiority of the Context Clustering architecture, achieving the highest AUC as well as the best ACC and F1 scores in single-view learning, indicating its balance in mammography tasks. $SV$ represents a single-view learning approach.
III-D2 Different Information Fusion Method
@@
The aim of this ablation study is to validate the effectiveness of our Tri level Information Fusion mechanism that combines these two types of local information with global information.
TABLE VI: Performance of different information fusion method on the Vindr-Mammo
@@
Global $0.783\pm 0.02$ 0.815 0.852
Global + Patch-based local $0.810\pm 0.02$ 0.890 0.891
Global + Feature-based local $0.806\pm 0.02$ 0.895 0.868
TIFF(ours) $\textbf{0.828}\pm 0.02$ 0.919 0.906
We can clearly found that focusing only on one type of local information does not produce the best results, but it is still significantly better than focusing only on Global Information. Adding Patch-based local information to Global Information will increase the AUC to 0.810, and adding Feature-based local information to Global Information will increase the AUC to 0.806. However, Tri level Information Fusion mechanism combining the three information achieved the best result with an AUC of 0.828.